In a low-energy transmission electron microscope (TEM), the typical average beam current is about -1.00 HA. If a material sample is exposed to this electron current for 17.8 min, how many electrons impact the material during that time? electrons
Answer #1

here,

current , I = 1 uA = 1 * 10^-6 A

time , t = 17.8 min = 1068 s

the charge passed , Q = I * t

Q = 1 * 10^-6 * 1068 C

Q = 1.068 * 10^-3 C

the number of electrons impacted , n = Q/e

n = 1.068 * 10^-6 /( 1.6 * 10^-19)

n = 6.68 * 10^12 electrons